# Hidden Breakthroughs in Language Model Training

### v2

@@ 4 Arithmetic language modeling
We validate our POLCA clustering method in a synthetic setting using an arithmetic addition task. Our clusters reflect categorical concepts within the data, even when those concepts are not discoverable by clustering directly on loss curves. Specifically, if we cluster on exact loss curves we recover digit positions, but if we cluster on POLCA curves we also recover the skill of “carrying” a digit.
@@ 4.1 Experiments
Our synthetic experiments use data from the arithmetic addition setting in Chen et al. [7], where the model is trained to compute the sum of two 3-digit numbers. As shown in Appendix Figure 7 and Chen et al. [7], the skills corresponding to the digits have different loss curves, so we will easily recover the digit skill categories by clustering exact loss curves. Unlike our source material, we also consider an additional skill: arithmetic carries to the output token (Figure 3). Digit-specific addition skills lead to clearly distinguishable loss curves, whereas carrying skills are not clear from the exact loss (Appendix Figure 8), but become clear on the decomposed gradient basis.
Setup details We train a 3-layer (9m parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048 [22]. We choose this model size to align with prior work Olsson et al. [24] and to maximize the granularity at which we can feasibly compute the POLCA values. For a validation set with 1250 data points and 5000 output tokens, we compute the loss and POLCA values for each token at intervals of 20 iterations throughout training. We choose the training steps between each POLCA computation to achieve as fine-grained analysis as possible without exploding the compute time. We compute the POLCA basis using the eigenvectors of the Hessian estimated using a 1250 data point sample of the training set as detailed in Algorithm 1. We compute one new basis vector every 200 iterations for a total of 50 basis vectors. We provide further ablations on the decomposition strategy and choice of POLCA basis in Appendix D and E.
@@
To ensure that we are investigating directions where the model is learning on average, we only consider the basis vectors for which the mean projected loss decreases. Then for each remaining basis vector, we remove all of the tokens for which the decomposed loss increases. We use HDBSCAN to cluster the remaining tokens, discarding the tokens it marks as outliers. Through this process, we find subpopulations of the data that have similar projected loss trajectories.
@@ 4.2 Results
In our clustering experiments on arithmetic addition skills, we first consider whether directional decomposition is necessary for identifying individual concepts. As a baseline, we therefore cluster tokens solely on their exact loss curves for successive timesteps, rather than estimating the loss decomposed along a low rank basis.
According to the HDBSCAN loss clustering results in Figure 3, we can recover—to a substantial degree—the digit skill by clustering only on the exact loss, likely because the digits have very different loss trajectories. However, as shown in Figure 3 and Appendix Table 3, we cannot recover clusters that are homogenous with respect to the carrying from the exact loss alone.
@@
Because exact loss clustering failed to recover the carrying skill, we instead cluster on each basis vector’s projected loss using the POLCA decomposition. The projected loss value (Equation 10) represents the cumulative loss change of attributed to movement along basis vector . By clustering the projected loss trajectories, we find that the top 3 basis vectors produce homogeneous clusters corresponding to both digit and carrying skills for each of the digits (Figure 4 and Appendix Figure 9), so POLCA clustering is able to recover subtler skills, like carrying, that are challenging to reconstruct from the exact loss curves alone.
Figure 4 shows the POLCA breakthrough clustering for the first three basis vectors. Along these directions, certain data subpopulations have changes in the projected loss (Figure 4(a) and 4(d)) that do not occur as visibly in their full loss curves (Figure 4(b), 4(e)). We conclude that arithmetic carries rely on breakthroughs along specific dimensions during training, but these breakthroughs may be elided in the exact loss curve.
@@ 5 English language modeling
5.1 Experiments
For the natural language modeling setting, we use the English Wikipedia dataset [35] from March 2022. We use the same POLCA setup as in the arithmetic addition setting. We train a 3-layer (40m parameter) transformer model with embedding dimension 512, 4 attention heads, and an MLP dimension of 2048 [22]. We compute the loss and POLCA values for each token on a validation set of 12600 output tokens. We analyze intermediate checkpoints at intervals of 200 iterations throughout training. We apply POLCA to the basis derived from the eigenvectors of the Hessian estimated using a 1000 data point sample of the training set as detailed in Algorithm 1 with . We compute a new basis vector every 750 iterations.
Similarly to the arithmetic setting, we only analyze the POLCA trajectories along the directions for which the decomposed loss decreases on average. We also discard any trajectories for which the decomposed per-token loss increases. After clustering the remaining token trajectories with HDBSCAN, we discard any marked as outliers before analysis.
@@ 5.2 Results
On language models, POLCA clustering again reveals hidden breakthroughs along each basis vector. To analyze the learned skills represented by each cluster, we look for syntactic and lexical patterns shared by the cluster data. After tagging tokens by part-of-speech (POS) using spacy, we label each cluster by computing its most common token and POS tag sequences. We confirm these automatically assigned cluster labels by examining the top 10 contexts closest to the medoid of the cluster.
We report a selection of clusters from specific basis vectors in Figure 6 (see Appendix H for more examples). Using POLCA clustering on projected loss trajectories for our basis vectors, we find token subpopulations that correspond to various grammatical constructions and have breakthroughs in the loss projected onto that basis. For instance, we show a cluster corresponding to predicting < to> and < from> after the first clause in a sentence in Figure 6a(6ii). We also observe clusters whose projected loss trajectories move in opposing directions along certain basis vectors. For instance, in Figure 6b(6i) and 6b(6ii), cluster 1 contains appositive noun phrases and cluster 2 has noun phrases with similar syntactic patterns that are not appositive (such as list items). These clusters visibly mirror each other’s movement in direction, if not in magnitude—decreases in loss are generally larger than the opposing cluster’s increases in loss which mirror them.
Figure 6 shows that despite their smooth exact loss curves, POLCA clusters have sudden changes in their decomposed loss curves at different points during training. Clusters from the exact loss curves, by contrast, do not reveal breakthroughs except very early in training (Appendix Figure 10). We conclude that POLCA reveals breakthroughs in the decomposed loss that are obscured in the exact loss. Through clustering, POLCA can explain how different skills are learned during language model training.
@@ 6 Conclusions
These are promising findings for meaningfully interpreting large models. By recovering breakthroughs in identifiable skills, we support the hypothesis that high-dimensional learning typically entails a series of phase transitions at various scales. When a phase transition appears in training, it suggests a naturally discrete category; the model either knows the concept or doesn’t know it, with little middle ground. Humans think in categorical concepts, so they are far more interpretable than the continuous data interpolations that appear in much of learning theory.
Limitations
We use small models so that we can have very fine-grained checkpoints for POLCA computation. Scaling this work to larger models may require using a basis that is less computationally expensive to compute than Hessian eigenvectors, but our results from Appendix E indicate that this is likely possible with limited impact on the cluster quality.
@@ Acknowledgements
Appendix A Derivation of approximate second-order term
@@
Note that the assumption above (of matching Hessian eigenvectors between data points and their aggregate) is unlikely to be correct. If this assumption is violated, then the scaling factor in the following second-order Taylor term will be minuscule on the datapoint level. In practice, we have found that the second-order term has limited impact at the datapoint level (see Appendix F), but we nonetheless use it to improve our approximation. Then we may approximate it as:
Appendix B Hyperparameters
@@
Appendix C Exact loss trajectories for the digit and carry skills
Appendix D Decomposition strategy comparison
We investigate whether POLCA is required for decomposing the loss. We test three strategies:
Loss: Exact loss
Fisher information: We approximate the empirical Fisher Information as as in [2]. For each basis vector , the Fisher Information projected onto is .
@@ Appendix D Decomposition strategy comparison
While using the projected gradient is less computationally intensive than computing the entire POLCA term, the results from Table 3 demonstrate that clustering on solely the Fisher information does not yield sufficiently homogeneous carry clusters.
| Decomposition strategy | Top vector skill homogeneities | ||
| 1 | 2 | 3 | |
| Loss | 0.514 | - | |
| Fisher information | 0.664 | 0.643 | 0.637 |
| POLCA | 0.973 | 0.946 | 0.903 |
Appendix E POLCA basis comparison

### v4

@@ 4 Arithmetic language modeling
We validate our POLCA clustering method in a synthetic setting using an arithmetic addition task. Our clusters reflect categorical concepts within the data, even when those concepts are not discoverable by clustering directly on loss curves. Specifically, if we cluster on exact loss curves, we recover digit positions—but if we instead cluster on POLCA curves, we also recover the skill of “carrying” a digit.
@@ 4.1 Experiments
Our synthetic experiments use data from the arithmetic addition setting in Chen et al. (2024b), where the model is trained to compute the sum of two 3-digit numbers. This setting has 4 skills corresponding to each of the digits in the output sum. Note that the digit in the 1000s place is always a <0> or <1> token since the two input summants are 3 digits long. As shown in Appendix Figure 7 and Chen et al. (2024b), the skills corresponding to the digits have different loss curves, so we will easily recover the digit skill categories by clustering exact loss curves.
Unlike our source material, we also consider an additional skill: arithmetic carries to the output token (Figure 3). This skill corresponds to the case where instead of simply adding the two tokens at the corresponding digit in the input, the model must also carry a from the previous digit. Digit-specific addition skills lead to clearly distinguishable exact loss curves, whereas carrying skills do not (Appendix Figure 8)—but carries will become clear on our decomposed gradient basis. We provide additional skill and labeling details in Appendix E.1.
Setup details We train a 3-layer (9 million parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048, following prior work (Olsson et al., 2022a). We study a validation set with 1250 data points and 5000 output tokens throughout training. We compute the loss and POLCA values for each token at each interval of 20 training steps. The POLCA basis uses the eigenvectors of the Hessian, estimated using a 1250 data point sample of the training set as detailed in Algorithm 1. We compute one new basis vector every 200 training steps, for a total of 50 basis vectors. We provide further ablations on the decomposition strategy and choice of POLCA basis in Appendices G and H. We train the model for steps, but trim the x-axis of the plots at to better display the breakthroughs.
@@
As described in Section 3.1, some of the top Hessian directions may represent directions of oscillation (and not learning) during training. To ensure that we are investigating directions where the model is learning on average, we only consider the basis vectors for which the mean projected loss decreases. Then for each remaining basis vector, we remove all of the tokens for which the decomposed loss increases, therefore retaining only tokens which rely on the vector during training. This removes 2360.8 out of 5000 tokens on average—suggesting that any given direction is irrelevant to learning many examples. We use HDBSCAN to cluster the remaining tokens, discarding the tokens it marks as outliers. Through this process, we find subpopulations of the data that have similar projected loss trajectories.
@@ 4.2 Results
| Exact loss | 0.514 | 0.0 | |||||||
| Change in exact loss | 0.524 | 0.0 | |||||||
| LCA (Lan et al., 2020) | 0.792 | 0.019 | |||||||
| POLCA | 0.973 | 0.355 |
@@
In our clustering experiments on arithmetic addition skills, we first consider whether decomposition is necessary for identifying conceptual skills. As a baseline, we therefore cluster tokens solely on their exact loss curves, rather than their decomposed loss curves. According to the exact loss clustering results in Figure 3, we can recover—to a substantial degree—the digit skill by clustering only on the exact loss, likely because the digits have very different loss trajectories. However, as shown in Figure 3 and Table 1, we cannot recover clusters that are homogenous with respect to the carrying skill from the exact loss alone.
@@
Because exact loss clustering failed to recover the carrying skill, we will recover it with a different clustering method. Instead of exact loss, we will cluster on each basis vector’s projected loss using the POLCA decomposition. The projected loss value (Equation 8) represents the cumulative loss change of attributed to movement along basis vector . By clustering the projected loss trajectories, we find that the top 2 basis vectors produce two easily-described and near-homogeneous clusters: one which contains examples of the 1000s place digit and one which contains examples of arithmetic carrying for all digits (Figure 4 Appendix Figure 9). We therefore determine that POLCA clustering recovers subtler skills—like carrying—that are challenging to reconstruct from the exact loss or parameter-aligned LCA curves alone (Table 1). In Appendix Table 9, we also use Equation 10 to compare the fraction of clusters with hidden breakthroughs past step (with ), where the mean exact loss plateaus (Appendix Figure 7), and find in Table 1 that POLCA discovers the highest fraction of clusters with hidden breakthroughs.
Figure 4 shows POLCA clustering for the first two basis vectors. Along these basis vectors, certain data subpopulations show changes in the projected loss (Figures 4(a) and 4(d)) that do not occur as visibly in their smoother exact loss curves (Figures 4(b) and 4(e)). We conclude that arithmetic carries rely on breakthroughs along specific dimensions during training, but these breakthroughs may be elided in the exact loss curve.
@@ 5 English language modeling
5.1 Experiments
For the natural language modeling setting, we use the English Wikipedia dataset (Wikimedia Foundation, 2022) from March 2022 to train a 3-layer (40m parameter) model. We use the same POLCA setup as in the arithmetic addition setting (see Appendix E.2 for details). As in the arithmetic setting, we only consider directions where the projected aggregate loss overall decreases to filter out directions of oscillation. We also discard the datapoint-specific POLCA trajectories along the directions where the token’s decomposed loss increases, which removes an average 6655.5 out of 12600 tokens per direction. After clustering the remaining token trajectories, we discard HDBSCAN-marked outliers.
Automatic labeling.
To analyze the concepts represented by each cluster, we look for syntactic and lexical patterns shared by the cluster data.111In principle, our method can uncover skills that are not describable through these templates, but templating allows automatic labeling. Our templating approach is similar to the automated explainer tool N2G (Foote et al., 2023), a popular ngram-based evaluation metric for unsupervised interpretability methods (Gao et al., 2024). Our automatic labels are based on the target token and its preceding trigram. To obtain automatic labels for a cluster, we compute the frequency of each POS tag (tagged by spacy) in it. We then automatically label each cluster with the smallest set of POS tags required to compose of the cluster’s 4-gram samples. For example, if over of the token instances in a cluster are preceded by a comma <,>, the cluster would be automatically labeled as 1 token after PUNCT. We consider basis vectors with at least one cluster with a simple label—one including at most two POS tags (not counting <PAD> tokens). Starting with 30 basis vectors, we remove 4 because the average decomposed loss increases. We find of the remaining basis vectors have at least one cluster with a simple label as defined. We refine these automatically assigned labels by examining the most frequent tokens in the cluster and the ten examples closest to the cluster medoid and manually selecting a label that is consistent with the automatic POS label and the ten closest examples to the medoid. See Appendix E.2 for further labeling details.
@@ 5.2 Results
On language models, POLCA clustering again reveals hidden breakthroughs along each basis vector. We show a selection of clusters from specific basis vectors in Figure 6 and others in Appendix K (Figures 12 and 14). Clustering projected loss trajectories along each basis vector, we find groups that correspond to various grammatical constructions. Some show apparent breakthroughs in their projected loss, like the cluster corresponding to predicting < to> and < from> after the first clause in a sentence (Figure 6a(6i)). We also observe clusters whose projected loss trajectories move in opposing directions along certain basis vectors. For instance, in Figure 6b, cluster 1 contains appositive noun phrases and cluster 2 has syntactically similar—but non-appositive—noun phrases (such as list items). These clusters visibly mirror each other’s movement—though the clustered decreases in loss are generally larger than their opposing cluster’s increases in loss.
Figure 6 shows that despite their smooth exact loss curves, POLCA clusters have sudden changes in their decomposed loss curves at different points during training. Clusters from the exact loss curves, by contrast, do not reveal breakthroughs except very early in training (Appendix Figure 10). We conclude that POLCA reveals breakthroughs in the decomposed loss that are obscured in the exact loss. Through clustering, POLCA can explain how different skills are learned during training.
@@ 6 Conclusions
These are promising findings for meaningfully interpreting large models. By recovering breakthroughs in identifiable skills, we support the hypothesis that high-dimensional learning typically entails a series of breakthroughs at various scales. When a breakthrough appears in training, it suggests a naturally discrete category; the model either knows the concept or doesn’t know it, with little middle ground. Humans think in categorical concepts, so these breakthroughs are far more interpretable than the continuous data interpolations often assumed in learning theory.
Reproducibility statement
We implement our models and experiments using open-source libraries and datasets. We provide detailed hyperparameters in Appendix D and a thorough explanation of the experimental setup in Section E. Our code is available at https://github.com/skangasl/POLCA.
Ethics statement
This work provides a method for better understanding the training dynamics of language models. The trained models may contain biases from the training datasets.
@@ Acknowledgements
Appendix A Limitations and future work
Our experiments are limited to small models. The two main challenges with scaling POLCA are the Hessian basis computation and the frequency of checkpoints used to sample the POLCA trajectories. Scaling this work to larger models may require using a basis that is less computationally expensive to compute than Hessian eigenvectors, but our results from Appendix H indicate that this is likely possible with limited impact on the cluster quality. The small scale of models that we use in our experiments allows for very high granularity of checkpoints used to compute both the basis and the POLCA trajectories. For larger models, this may be computationally infeasible and a lower checkpoint frequency may be needed, resulting in less signal for clustering the POLCA trajectories.
Our current experiments are limited to language models. However, in principle our approach is model-agnostic and can be applied to any deep neural network. Applying POLCA to other modalities is an exciting direction for future work.
The labeling approach that we use in the natural language setting relies on POS tagging. This labeling strategy allows for unsupervised, automatic identification of these syntactic skills and ensures strict interpretable labels. However, it fails to capture many human-interpretable language modeling skills. The discarded vectors may (and likely do) contain other interpretable skill clusters that are not found by automated labeling.
Appendix B Statement on use of large language models
We used generative AI for debugging and minor grammar edits when writing. The authors made all significant contributions to the research, analyses, and writing.
Appendix C Derivation of approximate second-order term
@@
Note that the assumption above (of matching Hessian eigenvectors between data points and their aggregate) is unlikely to be correct. If this assumption is violated, then the scaling factor in the following second-order Taylor term will be minuscule on the datapoint level. In practice, we have found that the second-order term has limited impact at the datapoint level (see Appendix I), but we nonetheless use it to improve our approximation. Then we may approximate it as:
Empirical Validation
We run a small-scale empirical study to test the accuracy of this estimation. For a sample of tokens and the first POLCA checkpoints and the top basis vectors in the arithmetic setting, we compute the root mean squared error (RMSE) between the approximated second-order term and the true second-order term averaged across tokens, checkpoints, and basis vectors. We find that our approximation has an RMSE of when compared to the ground truth second-order term, indicating that this approximation is close to the real value.
Appendix D Hyperparameters
@@
We selected the clustering hyperparameters to maximize empirical performance in the synthetic setting and used similar values relative to the number of tokens in the natural language experiments. We chose the POLCA hyperparameters to maximize the frequency of POLCA checkpoints and basis checkpoints within computational constraints.
@@
| Basis checkpoint interval (steps) | 200 |
| T | 50 |
| k | 1 |
| POLCA checkpoint interval (steps) | 5 |
| Hyperparameter | Value |
| Clustering algorithm | HDBSCAN |
| Minimum cluster size | 150 |
| Minimum samples | Number of tokens / 15 |
| Hyperparameter | Value |
@@
| Hyperparameter | Value |
| Basis checkpoint interval (steps) | 750 |
| T | 30 |
| k | 1 |
| POLCA checkpoint interval (steps) | 200 |
| Hyperparameter | Value |
| Clustering algorithm | HDBSCAN |
| Minimum cluster size | 300 |
| Minimum samples | Number of tokens / 15 |
Appendix E Experimental details
E.1 Arithmetic setting
Setup.
In the arithmetic experiments in Section 4, we train a 3-layer (9m parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048 (Nanda & Bloom, 2022). For a validation set with 1250 data points and 5000 output tokens, we compute the loss and POLCA values for each token at intervals of 20 steps throughout training. We compute the POLCA basis using the eigenvectors of the Hessian estimated using a 1250 data point sample of the training set. We compute a new basis vector every 200 steps for a total of 50 basis vectors.
Labeling details.
We automatically label each token with the ground truth value of the digit and carry skills using the definition of these two skills. We label the digit skill based on the position of the token in the output: the first token is 1000s, the second token is 100s, the third is 10s, and the fourth is 1s. We label the carry skill by computing the sum up to the next lowest digit place and determining whether it resulted in a carry to the current token. For instance, for an output in the 10s place, if the two inputs in the 1s place sum to 10 or higher, then it will be labeled with "carry", otherwise it will be labeled with "no carry".
E.2 Natural language setting
Setup.
For the natural language experiments in Section 5, we train on the English Wikipedia dataset (Wikimedia Foundation, 2022) from March 2022. We use the same POLCA setup as in the arithmetic addition setting. We train a 3-layer (40m parameter) transformer model with embedding dimension 512, 4 attention heads, and an MLP dimension of 2048 (Nanda & Bloom, 2022). We compute the loss and POLCA values for each token on a validation set of 12600 output tokens. We analyze intermediate checkpoints at intervals of 200 steps throughout training. We apply POLCA to the basis derived from the eigenvectors of the Hessian estimated using a 1000 data point sample of the training set as detailed in Algorithm 1 with . We compute a new basis vector every 750 steps.
Labeling details.
We label each token and the three tokens before it using spacy for part-of-speech (POS) tagging. This produces a sequence of four POS tags as the automatic POS label for each token. To label a given cluster, we compute the frequency (across the tokens in the cluster) of each POS tag at each index. For each index in the sequence, we then automatically label the cluster with the smallest set of POS tags at that index required to make up 70% of the cluster. We then filter out any labels that require more than 2 POS tags to describe the cluster. To generate the labels reported in Section 5, we manually refine the automatically generated label by looking at the top 10 contexts closest to the medoid of the cluster. Although these manually refined labels are challenging to verify automatically, we ensure that the contexts closest to the medoid follow the manual label and that the manual label follows the automatic label generated using the POS tags.
Appendix F Exact loss trajectories for the digit and carry skills
Appendix G Decomposition strategy comparison
We investigate whether POLCA is required for decomposing the loss. We expand on the results from Table 1 by showing the top three homogeneity scores for the carrying skill and adding empirical Fisher information and first order POLCA results. To compute the homogeneity score for a given basis vector, we take the maximum fraction of carries in any given cluster for that basis vector (or across all of the clusters for exact loss or change in exact loss). We then report the top three homogeneities across the full basis, as well as the recall and F1 score for the corresponding clusters. We note that we exclude HDBSCAN outliers from the recall and F1 computations. Importantly, we only consider clusters for which over of the tokens with the carry skill correspond to the s or s place, since carrying to the s place corresponds to simply predicting a or in the first position (see Figure 3 for reference) and is recovered with high homogeneity for all trajectory types except exact loss.
We compare carry skill homogeneity across the following trajectory types:
Loss: Exact loss trajectories.
Change in exact loss: We compute the change in exact loss by subtracting the loss at checkpoint from the loss at checkpoint for each timestep in the exact loss trajectory.
Fisher information: We approximate the empirical Fisher Information as as in Achille et al. (2017). For each basis vector , the Fisher Information projected onto is .
Loss Change Allocation (LCA): We compute the datapoint-wise LCA trajectories (Equation 1) projected onto the parameters that have the top highest magnitudes at the end of training.
First order POLCA: We calculate the POLCA trajectory without the second order term (Equation 2)
@@ Appendix G Decomposition strategy comparison
The results from Table 8 demonstrate that POLCA finds the most homogeneous clusters with respect to the carrying skill. Moreover, POLCA and first order POLCA have comparable F1 scores. Note that many low homogeneity clusters can have high recall because they are large and thus contain most of the carry tokens while having low homogeneity.
| Decomposition strategy | Cluster | |||
| Number | Homogeneity | Recall | F1 | |
| Loss | 1 | 0.514 | 0.771 | 0.617 |
| Change in exact loss | 1 | 0.524 | 0.958 | 0.678 |
| 1 | 0.664 | 0.947 | 0.781 | |
| Fisher information | 2 | 0.643 | 0.740 | 0.688 |
| 3 | 0.637 | 0.874 | 0.737 | |
| 1 | 0.792 | 0.772 | 0.782 | |
| Loss change allocation (LCA) (Lan et al., 2020) | 2 | 0.614 | 0.873 | 0.721 |
| 3 | 0.592 | 0.626 | 0.609 | |
| 1 | 0.948 | 0.767 | 0.848 | |
| First order POLCA | 2 | 0.928 | 0.769 | 0.841 |
| 3 | 0.887 | 0.751 | 0.813 | |
| 1 | 0.973 | 0.736 | 0.838 | |
| POLCA | 2 | 0.946 | 0.773 | 0.850 |
| 3 | 0.903 | 0.762 | 0.827 | |
We also compare the fraction of clusters with hidden breakthroughs for each type of trajectory. To compute whether or not a given cluster has a hidden breakthrough, we use Equation 10 with to identify breakthroughs past step where the exact loss plateaus. We find that POLCA produces the highest fraction of clusters with hidden breakthroughs. We hypothesize that this is mostly due to the basis construction, since the Fisher information and first order POLCA (both of which are computed using the same basis as POLCA) produce the next highest fraction of clusters with hidden breakthroughs.
| Decomposition strategy | Fraction of clusters with hidden breakthroughs |
| Loss | 0.0 |
| Change in exact loss | 0.0 |
| Fisher information | 0.284 |
| Loss change allocation (LCA) (Lan et al., 2020) | 0.019 |
| First order POLCA | 0.307 |
| POLCA | 0.355 |
Appendix H POLCA basis comparison